# A Selective Homomorphic Encryption Approach for Faster Privacy-Preserving Federated Learning

### v2

@@ VI-G Effects on Security
Table II compares partly encryption with full encryption across a series of security evaluations. The results confirm that partly encryption provides comparable protection, reinforcing its viability as a lightweight yet effective security mechanism.
VI-G1 Evaluating Privacy with MSSIM and VIFP
Model inversion attacks pose a threat to federated learning by reconstructing training data from model updates. To assess privacy resilience, we use Mean Structural Similarity Index (MSSIM) and Visual Information Fidelity in the Pixel domain (VIFP).
@@ VI-G1 Evaluating Privacy with MSSIM and VIFP
VI-H Cross-testing with Related Work
VI-H1 Comparison with MASKCRYPT
MASKCRYPT [19] applies a gradient-guided encryption mask to selectively encrypt critical model weights, reducing communication overhead by up to 4.15× compared to full model encryption. However, its gradient mask is not always optimized for selecting the most sensitive gradients, leading to lower initial security scores. MASKCRYPT requires multiple rounds to stabilize encryption, as reflected in VIFP scores, which only stabilize by round 5. This indicates an adaptation period where security is suboptimal.
In contrast, FAS stabilizes earlier by combining selective homomorphic encryption, differential noise, and bitwise scrambling, encrypting a fixed percentage of weights without gradient sensitivity analysis or pre-training. This approach reduces processing time by 90% compared to full homomorphic encryption while maintaining strong resistance to model inversion attacks. Unlike MASKCRYPT, which recalculates sensitivity masks each round, FAS eliminates extra computational costs and scales effectively across federated networks.
VI-H2 Comparison with FedML-HE
FedML-HE [18] identifies and encrypts critical weights based on gradient sensitivity through a pre-training phase. Our recreated FedML-HE model provides insight into its behavior, though results may differ from the original implementation. While FedML-HE offers strong privacy protection, its pre-training phase introduces significant time and resource costs. Each client generates an encryption mask independently, which, while avoiding direct mask sharing, can lead to inconsistencies in aggregated model accuracy.
Our FAS method eliminates the need for pre-training and mask aggregation by encrypting a fixed percentage of weights with differential noise and bitwise scrambling. At 10% encryption, FAS achieves a MSSIM score of 58%, comparable to FedML-HE’s 52%, but with significantly lower computational costs. The lack of pre-training and per-client mask generation allows FAS to maintain efficiency and scalability in federated learning environments.
@@ VI-H2 Comparison with FedML-HE
Figure 12 compares FAS, FedML-HE, and MASKCRYPT across MSSIM and VIFP scores. At 10% encryption, FedML-HE shows slightly higher stability (MSSIM 52%), while FAS achieves 58%, and MASKCRYPT falls in between at 55%. As encryption increases to 20%, all methods converge with nearly identical security performance. Despite slight advantages for FedML-HE in lower encryption levels, FAS provides comparable security at significantly lower computational costs. MASKCRYPT’s need for sensitivity mask recalculation each round makes it less efficient for real-time scenarios. These findings position FAS as an optimal choice for scalable, privacy-sensitive federated learning applications requiring low-latency performance.
TABLE III: Comparison of Encryption Techniques for Different Models
@@
Table III compares non-encrypted, partly encrypted, and overhead times for different models using FAS, FedML-HE, and MASKCRYPT on the Kidney dataset. The values are reconstructed based on methodologies described in the respective papers. FAS consistently has the lowest overhead across models, making it the most efficient. FedML-HE has the highest overhead due to pre-training, while MASKCRYPT offers a balance but initially has lower security, improving after multiple rounds. Overall, FAS is the best option for efficiency and scalability in federated learning.
VI-I Comparative Analysis of Encryption Techniques Across Datasets
@@
TABLE IV: Top 3 Cases Where FAS Outperforms Other Models
Dataset Model Compared To Enc. Improvement (%) Overhead Improvement (%)
@@
COVID EffNetB0 MASKCRYPT 36.36 60.61
Table IV presents the top three cases where the FAS encryption model demonstrates a higher efficiency compared to FEDML-HE and MASKCRYPT in terms of encryption time and overhead reduction.
@@
VI-J Handling Data Skew in Encryption-Based Techniques
In this section, we explore the impact of data skew on different encryption techniques, focusing on the robustness of the proposed FAS method compared to existing approaches such as MASKCRYPT and Fedml-HE.
VI-J1 Challenges with Mask-Based Techniques
Both MASKCRYPT and Fedml-HE rely on sensitive masks to select important gradients for encryption. These masks are designed to prioritize gradients that have a significant impact on the model’s accuracy. However, in scenarios with skewed data distributions, the importance of gradients becomes difficult to determine accurately. As a result, the encryption decisions made by these methods often degrade to levels resembling random encryption. This reduces their effectiveness, as observed in the MSSIM and VIFP score comparisons.
VI-J2 Robustness of the FAS Technique
In contrast, the proposed FAS technique, which operates as a random encryption method, is unaffected by data skew. Unlike mask-based techniques, FAS does not depend on the model’s accuracy or gradient importance for its encryption process. Instead, it employs a combination of encryption, scrambling, and noise addition. These components ensure that the method maintains consistent performance irrespective of the underlying data distribution. The robustness of this approach makes it particularly suitable for scenarios where data skew is prevalent, such as federated learning environments with heterogeneous clients.
VI-J3 Experimental Results and Comparisons
Dataset Method MSSIM (Skew/Normal) VIFP (Skew/Normal)
@@ VI-J3 Experimental Results and Comparisons
TABLE V: MSSIM and VIFP Scores for Skew and Normal Data across Datasets
Table V demonstrates that the proposed FAS technique, leveraging encryption, scrambling, and noise, exhibits superior robustness to data skew across all evaluated datasets—Kidney, Lung, COVID, and Diabetic Retinopathy—when compared to existing gradient-based encryption methods. This characteristic ensures that FAS maintains its performance and security advantages even in diverse, real-world scenarios with varying data distributions. Experimental results using the EfficientNetB0 model across these datasets show that the accuracy of FAS is minimally affected by data skew, while other methods experience significant performance degradation under similar conditions. These findings highlight the adaptability and reliability of the FAS technique in privacy-preserving machine learning applications.
  • For MASKCRYPT and Fedml-HE, the MSSIM and VIFP scores under skewed data are significantl degraded, closely resembling random encryption.
  • In contrast, the FAS technique maintains reduced degregation across all conditions, as its random encryption approach is independent of the data distribution.
@@
VII Conclusion
This paper evaluated privacy-preserving techniques in federated learning, focusing on different datasets and comparing differential privacy, homomorphic encryption, and our custom FAS approach.
Full homomorphic encryption offers the highest data protection but incurs significant computational costs, making it suitable only for scenarios prioritizing confidentiality over performance. Differential privacy provides lightweight privacy with minimal impact on computation, ideal for moderate security requirements. Our FAS method strikes a balance by achieving security comparable to full encryption while significantly reducing training time and overhead, making it efficient for large-scale federated learning and resource-constrained environments.

### v4

@@ 6.7 Effects on Security
Table 2 compares FAS with FHE across a series of security evaluations. The results confirm that FAS provides comparable protection, reinforcing its viability as a lightweight yet effective security mechanism.
6.7.1 Evaluating Privacy with MSSIM and VIFP
Model inversion attacks pose a threat to FL by reconstructing training data from model updates. To assess privacy resilience, we use Mean Structural Similarity Index (MSSIM) and Visual Information Fidelity in the Pixel domain (VIFP).
@@ 6.7.1 Evaluating Privacy with MSSIM and VIFP
6.8 Cross-testing with Related Work
6.8.1 Comparison with MASKCRYPT
MASKCRYPT 10506637 applies a gradient-guided encryption mask to selectively encrypt critical model weights, reducing communication overhead by up to 4.15× compared to full model encryption. However, its gradient mask is not always optimized for selecting the most sensitive gradients, leading to lower initial security scores. MASKCRYPT requires multiple rounds to stabilize encryption, as reflected in VIFP scores, which only stabilize by round 5. This indicates an adaptation period where security is suboptimal.
In contrast, FAS stabilizes earlier by combining selective homomorphic encryption, differential noise, and bitwise scrambling, encrypting a fixed percentage of weights without gradient sensitivity analysis or pre-training. This approach reduces processing time by 90% compared to FHE while maintaining strong resistance to model inversion attacks. Unlike MASKCRYPT, which recalculates sensitivity masks each round, FAS eliminates extra computational costs and scales effectively across federated networks.
6.8.2 Comparison with FedML-HE
FedML-HE jin2024fedmlheefficienthomomorphicencryptionbasedprivacypreserving identifies and encrypts critical weights based on gradient sensitivity through a pre-training phase. Our recreated FedML-HE model provides insight into its behavior, though results may differ from the original implementation. While FedML-HE offers strong privacy protection, its pre-training phase introduces significant time and resource costs. Each client generates an encryption mask independently, which, while avoiding direct mask sharing, can lead to inconsistencies in aggregated model accuracy.
Our FAS method eliminates the need for pre-training and mask aggregation by encrypting a fixed percentage of weights with differential noise and bitwise scrambling. At 10% encryption, FAS achieves a MSSIM score of 58%, comparable to FedML-HE’s 52%, but with significantly lower computational costs. The lack of pre-training and per-client mask generation allows FAS to maintain efficiency and scalability in FL environments.
@@ 6.8.2 Comparison with FedML-HE
Figure 12 compares FAS, FedML-HE, and MASKCRYPT across MSSIM and VIFP scores. At 10% encryption, FedML-HE shows slightly higher stability (MSSIM 52%), while FAS achieves 58%, and MASKCRYPT falls in between at 55%. As encryption increases to 20%, all methods converge with nearly identical security performance. Despite slight advantages for FedML-HE in lower encryption levels, FAS provides comparable security at significantly lower computational costs. MASKCRYPT’s need for sensitivity mask recalculation each round makes it less efficient for real-time scenarios. These findings position FAS as an optimal choice for scalable, privacy-sensitive FL applications requiring low-latency performance.
Table 3: Comparison of Encryption Techniques for Different Models
@@
Table 3 compares non-encrypted, partly encrypted, and overhead times for different models using FAS, FedML-HE, and MASKCRYPT on the Kidney dataset. The values are reconstructed based on methodologies described in the respective papers. FAS consistently has the lowest overhead across models, making it the most efficient. FedML-HE has the highest overhead due to pre-training, while MASKCRYPT offers a balance but initially has lower security, improving after multiple rounds. Overall, FAS is the best option for efficiency and scalability in FL.
6.9 Comparative Analysis of Encryption Techniques Across Datasets
@@
Table 4: Top 3 Cases Where FAS Outperforms Other Models
Dataset Model Compared Enc. Impr. (%) Overhead Impr. (%)
@@
COVID EffNetB0 MASK-CRYPT 36.36 60.61
Table 4 presents the top three cases where the FAS encryption model demonstrates a higher efficiency compared to FEDML-HE and MASKCRYPT in terms of encryption time and overhead reduction.
@@
6.10 Handling Data Skew in Encryption-Based Techniques
This section investigates how data skew affects different encryption techniques, with an emphasis on evaluating the robustness of the proposed FAS approach in comparison to existing methods like MASKCRYPT and Fedml-HE.
6.10.1 Challenges with Mask-Based Techniques
Both MASKCRYPT and Fedml-HE rely on sensitive masks to select important gradients for encryption. that strongly affect how well the model performs. However, in scenarios with skewed data distributions, the importance of gradients becomes difficult to determine accurately. As a result, the encryption decisions made by these methods often degrade to levels resembling random encryption. This reduces their effectiveness, as observed in the MSSIM and VIFP score comparisons.
6.10.2 Robustness of the FAS Technique
In contrast, the proposed FAS technique, which operates as a random encryption method, is unaffected by data skew. Unlike mask-based techniques, FAS does not depend on the model’s accuracy or gradient importance for its encryption process. Instead, it employs a combination of encryption, scrambling, and noise addition. These components ensure that the method maintains consistent performance irrespective of the underlying data distribution. The robustness of this approach makes it particularly suitable for scenarios where data skew is prevalent, such as FL environments with heterogeneous clients.
6.10.3 Experimental Results and Comparisons
Dataset Method MSSIM (Skew/
Normal)
VIFP (Skew/
Normal)
@@ 6.10.3 Experimental Results and Comparisons
Table 5: MSSIM and VIFP Scores for Skew and Normal Data across Datasets
Table 5 demonstrates that the proposed FAS technique, leveraging encryption, scrambling, and noise, exhibits superior robustness to data skew across all evaluated datasets—Kidney, Lung, COVID, and Diabetic Retinopathy—when compared to existing gradient-based encryption methods. This characteristic ensures that FAS maintains its performance and security advantages even in diverse, real-world scenarios with varying data distributions. Experimental results using the EfficientNetB0 model across these datasets show that the accuracy of FAS is minimally affected by data skew, while other methods experience significant performance degradation under similar conditions. These findings highlight the adaptability and reliability of the FAS technique in privacy-preserving machine learning applications.
  • For MASKCRYPT and Fedml-HE, the MSSIM and VIFP scores under skewed data are significantly degraded, closely resembling random encryption.
  • In contrast, the FAS technique maintains reduced degragation across all conditions, as its random encryption approach is independent of the data distribution.
@@
7 Conclusion
This paper evaluated privacy-preserving techniques in FL, focusing on different datasets and comparing differential privacy, HE, and our custom FAS approach.
FHE offers the highest data protection but incurs significant computational costs, making it suitable only for scenarios prioritizing confidentiality over performance. Differential privacy provides lightweight privacy with minimal impact on computation, ideal for moderate security requirements. Our FAS method strikes a balance by achieving security comparable to full encryption while significantly reducing training time and overhead, making it efficient for large-scale FL and resource-constrained environments.